Hey i have a problem i can't solve this question can you help me. it's solving w.
w/w-1=3/w+1+1

Answer by stanbon(75887)

w/w-1=3/w+1+1
w/(w-1) = 3/(w+1) + 1
Multiply thru by (w-1)(w+1) to get:
w(w+1) = 3(w-1) + w^2-1
w^2 + w = 3w - 3 + w^2 -1
-2w = -4
w = 2
=============
Cheers,
Stan H.